# Generated Data with Fake Privacy: Hidden Dangers of Fine-tuning Large Language Models on Generated Data

### v1

@@ 2.4 Datasets
We evaluate the privacy risks on Pythia’s training dataset, the Pile. Especially, we use its Enron email subset and FreeLaw subset. Details are listed below.
2.4.1 Pile
Developed by EleutherAI [16], The Pile involves 800GB data from different sources including internet forums, video subtitles, and academic texts. The Pile has been used for various model’s pre-training such as GPT-Neo and Pythia. It consists of 22 smaller datasets that include Enron and Freelaw corpora.
2.4.2 Enron
Enron corpus [21] is a Pile subset containing different email conversations. We use a preprocessed version of the dataset shared by [33] which consists of 3330 samples. For each sample, they split the original sample in Enron into the following columns:
@@ 2.4.2 Enron
    Continuation. The second part of each selected conversation that completes the logical flow introduced by the prompt. LLM’s generation is compared with this column in terms of language, semantic similarity, and coherence.
@@
2.4.3 FreeLaw
FreeLaw is a free and published dataset related to the legal domain. It is a subset of The Pile is obtained from the CourtListener [4] project. CourtListener includes a large number of legal opinions from federal and state courts. It consists of a lot of modalities of legal proceedings, including dockets, bibliographic information on judges, and other metadata. Following Pile’s setting, we only focus on court opinions due to an abundance of full-text entries.
3 Privacy Risks on Fine-Tuning with Unstructured Generated Data.
@@ 2.4.3 FreeLaw
3.1 Experiment Settings
Figure 1: An overview of the privacy evaluation procedure for plain fine-tuning without generated data.
Since Pythia’s training data is open-sourced and allows for easy evaluation of privacy leakage, we chose Pythia as the base model for our experiments. To evaluate the potential risks, we first use Pythia-12b [7] as a generator to generate an email-related dataset. Then we fine-tune Pythia-410m, 1.4b, and 2.8b models on these data and evaluate the privacy risks with PII attacks following the pipeline as drawn in Figure 1. The details for the data generation, model fine-tuning, and evaluation are listed below.
@@ 3.1.1 Dataset Generation
We adopt the first $2220$ rows of the processed version of the processed Enron email dataset provided by [33] for data generation, denoted as the "seed" set. The seed split’s prompt column is used to generate alternatives for the continuation.
For the generation, we have used zero-shot prompting without any instruction template, that is the model predicts the next tokens based on the raw text to extend it. This is because the pythia-12b model is not fine-tuned on an instruction template.
@@
The used top-k value is $100$, and the temperature is $0.75$. The temperature is deliberately chosen high to introduce randomness and diversity into the generations. This approach is particularly relevant as it has been observed that such generations may include random PIIs, such as emails and company names. Previous research has noted that memorized data can be leaked through specific prompts or prefixes [19, 11]. However, PIIs may not always appear as the highest-likelihood output, meaning they might not be revealed in a greedy decoding setting. By using a higher temperature, we increase the likelihood of generating memorized data, enabling a more thorough assessment of data privacy risks.
@@ 3.1.2 Model Fine-Tuning
After obtaining the generated data related to Enron emails, we do 4-bit QLoRA [13] fine-tuning for Pythia-410M, 1.4b, and 2.8b with Adam optimizer [20] for $4$ epochs with $64$ rank, $1/4$ scale factor, $0.05$ LoRA dropout rate, and batch size of $8$. We searched the learning rate from the $42e-$ 4 to $85e-$ 8 to obtain the fine-tuned models with the most privacy risks. After merging the adapters with the appropriate base model and converting the parameters into 16-bit, we can get the fine-tuned Pythia models with enhanced ability in the email domain for inference.
@@ 3.2 Model Evaluation
Validation Data. To test the utility and privacy risks related to the fine-tuned models, we choose the first $1000$ rows of the Enron email dataset provided by [33] to form the validation data.
Utility For utility evaluation, we measure mean perplexity, a widely recognized metric based on negative log-likelihood, on the validation data to compare the fine-tuned models against their base counterparts. Conducting utility evaluation is crucial to simulate realistic use cases.
Privacy For the privacy evaluation, we first process the validation data into four templates, labeled A, B, C, and D, following Wang et al.[33]’s setting as follows,
@@
{name} and {email} are placeholders here. Following the 5-shot attack setting, we concatenate five samples of filled ’name’ and ’email address’ pairs within a selected template, preceeding one sample that contains only the ’name’ chosen from the evaluation set. These concatenated sentences are then formed as the inputs of the privacy validation data for PII attacks and use the ground truth email of the final name as the targets. After feeding the inputs of former defined privacy validation data, we then evaluate the privacy leakage by implementing the fuzzy string matching [27] method on generated emails and the ground truth. It is a commonly used metric to check if a string has a clear match with a given string based on the Levenshtein distance. If the similarity score is greater or equal to $80$, we will judge the sensitive email information is leaked by the evaluated model following Neel et al.’s [17] setting. The overall procedure is depicted in Figure 1.
Table 1: Number of successful extractions for different Pythia models and their perplexity across the four templates.
@@ 3.3.1 Worst Cases on Privacy Leakage
Following the above fine-tuning strategy and evaluation methods, we get the number of successful PII extractions of different models. Firstly, we list the highest number of successful extractions of different models across the four templates in Table 1 with their perplexities on the validation split of Enron email.
From Table 1, one can see that the model utilities evaluated by the perplexity are improved after fine-tuning, demonstrating the generated data successfully improves the models’ understanding of the email domain. However, there are more than $20\%$ improvements for the successful PII extractions after fine-tuning the generated data on models with different sizes, especially for the Pythia with 410M parameters. Such improvements demonstrate that fine-tuning generated data can lead to more serious privacy leakage on data related to the same domain although it can also effectively improve LLM’s knowledge on the related domain.
Furthermore, we observe that the number of successful PII extractions increases with model size in both the base and fine-tuned models, aligning with previous research [26]. Such a trend can be attributed to the superior representational capacity of larger models, which allows them to more easily memorize training data, as discussed in Tirumala et al.’s work [32]. Consequently, larger models not only exhibit improved performance but also present greater risks of successful PII extractions after fine-tuning with generated data. This highlights significant privacy concerns, especially as the development of larger LLMs continues to gain traction.
@@ 3.3.2 Results for Different Attack Templates
Except for reporting the highest number of successful PII extractions across different templates, we also draw different models’ PII behavior on each specific extraction template in Figure 2.
Figure 2: The number of successful extractions with different templates for the fine-tuned model (denoted as Generated data) and the base model.
Figure 2 demonstrates that the Pythia models consistently become more susceptible to extracting sensitive emails after fine-tuning with generated data, regardless of the evaluation template used. Especially for templates A, B, and C, most models’ successful PII extractions increase significantly. As for template D, we find the differences between the fine-tuned models’ successful PII extractions and the base models’ are less significant. A possible explanation is that template D contains more email-specific information, such as irregular characters in "Original Messages", which enhances the model’s memorization of email patterns, leading to accurate extractions even in the base models. As noted by previous research  [11, 19], special prompt prefixes with appropriate context increase data extraction success in base models. Templates that include more human language tend to be more successful on par with model size, i.e., the capability of natural language understanding.
@@
Figure 3: An example case for the PII attacks.
We draw an example to show the fine-tuned Pythia model can release more private information in Figure 3. From the figure, one can see before the fine-tuning, the base model can only infer the email address by simply rephrasing the given name with a randomly generated domain. However, one can see that the term "jdoe" and "energy" exist in the fine-tuned models’ response. These specific terms demonstrate that the fine-tuned model invokes more memorization of the pertaining data and causes potential privacy risks.
@@
Figure 4: The model’s perplexities and the number of successful extractions with respect to different learning rates used for fine-tuning. The horizontal line here denotes the perplexity of the base model. The x-axis here denotes the learning rate.
@@ 3.4 Discussion
3.4.1 Reasons for the Privacy Risks
To find the possible reason for the increased privacy risks, we first show a sample from the generated dataset in 5. From the figure, one can see that the structure of the generated data and test data are similar. Specifically, we noticed that the generated data included a lot of name image pairs. We also find that the total generated data consists of over $60,000$ name-email pairs for around $2000$ generated samples. Such structure similarities may revoke LLMs’ memorization of name and email pairs and lead to privacy risks, as LLMs feed too many name-email pairs during the fine-tuning.
Figure 5: An example of the generated data for fine-tuning.
Apart from the structure similarity, we also notice that the semantic similarities of the generated data are also similar. The semantic similarity score between the generated data and the original datasets is over $0.7$ evaluated by the Sentence Transformer. From Figure 5, we can also see many specific name-email relations are similar in both generated and test data. For example, the email’s local name can be formed by inserting a "_" or "." between the first name and the second name, or just concatenating the first character in the first name with the last name to form an email address like dhansen for Don Hansen. Such similar relationships exist in both the generated and the original pre-training data can recall LLM’s memory of the relationship between the name and emails.
3.4.2 Learning Rates Impacts on the Privacy Risks
As illustrated in Tirumala et al.’s[32] work, the learning rate is a key factor in LLM’s memorization and the model’s final performance. Therefore, we further explore the learning rate’s impacts on both the model’s utility and the extraction success rate in Figure 4. From the curves in Figure 4, one can see that the utility and extraction success rates show a similar trend with the learning rate changes. With the increment of the learning rate for fine-tuning, the perplexities after fine-tuning first decrease a lot but will also increase when the learning rate becomes too large. The best learning rate for better utility is around $10^{-6}$ to $5\times 10^{-5}$ for different models. As for the PII extractions, there are no clear trends for different models with respect to different learning rates. However, we can still observe one consistent pattern the successful PII extractions for models fine-tuned with smaller learning rates are not too high. The possible reason is that the slightly larger learning rate will help LLMs memorize better and lead to higher privacy risks, consistent with the findings in Tirumala et al.’s[32] work. Therefore, we recommend using the smaller learning for fine-tuning, like around $10^{-6}$, to alleviate the privacy risks while improving the utilities on the target domain.
3.5 Take-away Messages
In this section, we explore the privacy risks posed to LLMs on their pre-training datasets after fine-tuning with generated instructional data. Using the Enron email dataset as a case study, we fine-tune Pythia models of varying sizes (410M, 1.4B, 2.8B) with email data generated by a Pythia 12B model. We then assess the privacy risks by performing PII (Personally Identifiable Information) attacks on the Enron dataset, which is related to the pre-training data. The results reveal that, after fine-tuning, the Pythia models are able to extract over $20\%$ more successful PII data compared to the base model. This finding indicates that fine-tuning with generated data can heighten the model’s privacy risks concerning the pre-training dataset. Further analysis suggests that these risks may arise because the generated data includes structures or relationships similar to those in the pre-training dataset, triggering the LLMs to recall stored information. We also find that using a lower learning rate can mitigate these privacy risks by reducing the number of successful PII extractions, offering a potential solution for alleviation.

### v2

@@ 2.5 Datasets
We evaluate the privacy risks on Pythia’s training dataset, the Pile. Especially, we use its Enron subset for plain-text fine-tuning and FreeLaw subset for instruction-tuning.
Pile: The Pile [24] involves 800GB data from various sources including Internet forums, video subtitles, and academic texts. The Pile has been used for various model’s pre-training such as GPT-Neo and Pythia. It consists of 22 smaller datasets including Enron and Freelaw corpora.
Enron: Enron corpus [31] is a Pile subset containing different email conversations. We use a preprocessed version of the dataset shared by [51] which consists of 3330 samples. For each sample, the original sample in Enron is split into the following columns:
@@
    Continuation. The second part of each selected conversation completes the logical flow introduced by the prompt. LLM’s generation is compared with this column in terms of language, semantic similarity, and coherence.
@@
Psedonymized Enron: We create an extended version of the Enron dataset, referred to as the Pseudonymized Enron dataset, for experimental purposes where the pretraining data does not overlap with the generated fine-tuning data. In this dataset, the original names and email addresses from Enron were replaced with synthetic data generated using the Faker library [23]. Note that due to the randomness and the limited selection of names and emails in this library, some pseudonymous names and email addresses might appear multiple times. Names are identified and replaced using regular expressions to match capitalized words, while email addresses are detected by identifying patterns containing a domain following the "@" symbol.
FreeLaw: FreeLaw is an open-source dataset related to the legal domain. It is a subset of The Pile that is obtained from the CourtListener [4] project. CourtListener includes a large number of legal opinions from federal and state courts. It consists of numerous modalities of legal proceedings, including dockets, bibliographic information on judges, etc. Following Pile’s setting, we only focus on court opinions due to an abundance of full-text entries.
3 Privacy Risks on Fine-Tuning with Unstructured Generated Data
@@
3.1 Experimental Setting
Figure 1: An overview of the privacy evaluation procedure for plain fine-tuning with generated data. The generator LLM creates a complete fine-tuning corpus, introducing PII leakage risks. For instance, if an email is addressed to ’John,’ the generator might infer additional context, such as a vacation in Paris. When the complete context, including the initial email information, is used for fine-tuning, it heightens the risk of PII leakage.
Since Pythia’s training data is open-source and allows for easy evaluation of privacy leakage, we chose Pythia as the base model for our experiments. To evaluate the potential risks, we first use Pythia-12b [11] as a generator to generate an email-related dataset. Then we fine-tune Pythia-410m, 1.4b, and 2.8b models on these data and evaluate the privacy risks with PII attacks following the pipeline as drawn in Figure 1. The details for the data generation, model fine-tuning, and evaluation are listed below.
@@ 3.1.1 Dataset Generation
We adopt the first $2220$ rows of the processed version of the Enron email dataset provided by [51] for data generation, denoted as the "seed" split. The seed split’s prompt column (see Section 2.5) is used to generate alternatives for the continuation. For the generation, we use zero-shot prompting without any instruction template, that is the model predicts the next tokens based on the raw text to extend it. This is because the Pythia-12b model is not fine-tuned on an instruction template.
@@
The used top-k value is $100$, meaning only the most probably 100 next tokens is considered for generation. The temperature is chosen as $0.75$ to introduce significant randomness and diversify the generations. This approach is particularly relevant as it has been observed that such generations may include random PIIs, such as emails and company names. Previous research has noted that memorized data can be leaked through specific prompts or prefixes [28, 16]. However, PIIs may not always appear as the highest-likelihood output, meaning they might not be revealed in a greedy decoding setting. Using more diverse generations, we enable a more thorough assessment of privacy risks.
@@ 3.1.2 Model Fine-Tuning
After obtaining the generated data related to Enron emails, we do 4-bit QLoRA [19] fine-tuning for Pythia-410M, 1.4b, and 2.8b with Adam optimizer [30] for $4$ epochs with $64$ rank, $1/4$ scale factor, $0.05$ LoRA dropout rate, and batch size of $8$. We explored learning rates ranging from $42e-$ 4 to $85e-$ 8 to obtain the fine-tuned models with the most privacy risks. By merging the adapters with the appropriate base model and converting the parameters to 16-bit precision, we obtain fine-tuned Pythia models optimized for enhanced performance in the email domain during inference.
@@ 3.2 Model Evaluation
Validation Data. To test the utility and privacy risks related to the fine-tuned models, we choose the first $1000$ rows of the Enron email dataset [51] to form the validation data.
Utility. For utility evaluation, we measure mean perplexity, a widely recognized metric based on negative log-likelihood, on the validation data to compare the fine-tuned models against their base counterparts. Conducting utility evaluation is crucial to simulate realistic use cases.
Privacy. For the privacy evaluation, we first process the validation data into four templates, labeled A, B, C, and D, following Wang et al.[51]’s setting as follows:
@@
{name} and {email} are placeholders here. Following the 5-shot attack setting, we concatenate five samples of filled ’name’ and ’email address’ pairs within a selected template, preceding one sample that contains only the ’name’ chosen from the evaluation set. These concatenated sentences serve as input for the privacy validation dataset used to evaluate PII attacks, with the ground truth email associated with the final name designated as the target. After feeding the inputs of former defined privacy validation data, we then evaluate the privacy leakage by implementing the fuzzy string matching [44] method on generated emails and the ground truth. It is a commonly used metric to check if a string has a clear match with a given string based on the Levenshtein distance. Additionally, we explore for the best learning rate that achieves the highest attack success. If the similarity score is greater or equal to $80$, we will judge the sensitive email information is leaked by the evaluated model following Neel et al.’s [26] setting. The overall procedure is summarized in Figure 1.
@@
Table 1: The number of successful extractions for different Pythia models and their perplexity across the four templates.
@@ 3.3 Results
In this section, we provide the results on our experiments to showcase the privacy risks of fine-tuning with unstructured data. We first present the worst cases and then provide results for different attack templates. Finally, we discuss the reasons and learning rate impact on the privacy risks.
@@ 3.3.1 Worst Cases on Privacy Leakage
Following the fine-tuning strategy and evaluation methods explained in Section 3.1, we get the number of successful PII extractions of various models. Firstly, we list the highest number of successful extractions of different models across the four templates in Table 1 with their perplexities on the validation split of Enron email. For the Facebook OPT experiments, we omit these results from the table as the model size remains constant. The perplexities of these experiments are as follows: baseline model achieves 9.15, fine-tuned with DoRA achieves 8.25, and fine-tuned with LoRA achieves 8.18. We note that the impact of perplexity reduction depends on factors such as the baseline, dataset complexity, and specific application. To provide context, we compare our perplexity values against the base model. Our observations show that model utility, as indicated by perplexity, improves after fine-tuning. While comparable results are not available for the same dataset, literature suggests that a 3-10% reduction in perplexity is typically considered a meaningful improvement [41, 8]. However, we observe over a $20\%$ improvement in successful PII extractions after fine-tuning the models with the generated data, particularly for the Pythia model with 410M parameters. Such improvements demonstrate that fine-tuning with generated data can lead to more serious privacy leakage on data related to the same domain although it can also effectively improve LLM’s knowledge on the related domain.
Furthermore, we observe that the number of successful PII extractions increases with model size in both the base and fine-tuned models, consistent with findings from previous research [40]. This trend can be attributed to the enhanced representational capacity of larger models, which enables them to memorize training data more effectively, as highlighted in [49]. Consequently, larger models not only exhibit improved performance but also present greater risks of successful PII extractions after fine-tuning with generated data. This highlights significant privacy concerns, especially as the development of larger LLMs continues to gain traction.
@@ 3.3.2 Results for Different Attack Templates
Pythia Model Results. In addition to reporting the highest number of successful PII extractions across various templates, we also analyze the PII extraction behavior of different models for each specific template, as illustrated in Figure 2.
Figure 2: The number of successful extractions with various templates for the fine-tuned model (denoted as Generated data) and the base model on Pythia models.
Figure 3: An example case for the PII attacks.
We observe that the Pythia models consistently become more susceptible to extracting sensitive emails after fine-tuning with generated data, regardless of the evaluation template used. Especially for templates A, B, and C, we observe substantial elevation in extraction success for all models. For template D, we observe that the differences in successful PII extractions between the fine-tuned models and the base models are less pronounced. A possible explanation is that template D includes more email-specific information, such as irregular characters in "Original Messages," which aids the models in memorizing email patterns. This familiarity enables even the base models to perform accurate extractions. As noted by previous research  [16, 28], special prompt prefixes with appropriate context increase data extraction success in base models. Templates that include more human language tend to be more successful on par with model size, i.e., the capability of natural language understanding.
@@
Figure 4: The number of successful extractions with various templates for the fine-tuned model (denoted as Generated data) and the base model for Pythia using LoRA and DoRA as PEFT methods. All names and email addresses in the Enron dataset have been replaced with pseudonyms.
Figure 5: The number of successful extractions with various templates for the fine-tuned model (denoted as Generated data) and the base model for Facebook OPT 2.7b using LoRA and DoRA as PEFT methods.
Figure 6: An example of the generated data for fine-tuning.
@@
Figure 7: The model’s perplexities and the number of successful extractions with respect to various learning rates for fine-tuning the Pythia model. The horizontal line here denotes the perplexity of the base model. The x-axis here denotes the learning rate.
We draw an example to show the fine-tuned Pythia model can release more private information in Figure 3. As shown in the figure, prior to fine-tuning, the base model could only infer the email address by rephrasing the given name and appending a randomly generated domain. However, the term "jdoe" and "energy" exist in the fine-tuned models’ response. These specific terms demonstrate that the fine-tuned model invokes more memorization of the pertaining data and causes potential privacy risks.
We also evaluate the Pythia 1.4b model using the Pseudonymized Enron dataset, which exhibits minimal overlap with the sensitive data targeted for inference in Pythia’s pretraining dataset. Consequently, this represents a challenging experimental setup in the Pythia experiments. As shown in Figure 4, the model’s performance across most templates either matched or exceeded its previous results. Among the templates, Template A demonstrates the best performance. This behavior can potentially be attributed to the characteristics of the dataset and the model’s training corpus. A significant portion of Pythia’s training data consists of non-email content, which may align well with Template A’s similarity to non-email patterns and the repeated occurrences of sensitive data within Pseudonymized Enron. Overall, the LoRA PEFT method demonstrates slightly better attack success compared to the DoRA PEFT method. This difference may be due to LoRA’s superior performance in tasks requiring higher memorization capabilities, where DoRA slightly lags behind [29]. However, despite these distinctions, the overall performance gap between the two methods remains minimal.
We note here that templates B and C were particularly challenging for these experiments because the patterns used in these templates are almost exclusive to email data. For instance, Template C includes patterns such as [mailto: {email}], which closely resemble the structure of email headers, as seen in examples like [mailto: example@domain.com]. Similarly, Template B involves email-specific constructs such as name: {name}, email: {email}, mirroring common formats in structured email-related data, as illustrated by name: John Doe, email: doejohn@haymail.com. These examples underscore the difficulty of these templates in the Pseudonymized Enron setting, where the model has not encountered this type of data, in contrast to the original Enron setting discussed earlier.
We conclude that templates have a substantial impact on PII extraction success depending on the tasks present in the pretraining and fine-tuning data. Across all cases, Template B consistently demonstrates the lowest PII extraction success. Template C, with its precise and structured format, performs best when there is overlap between the pretraining and fine-tuning tasks, as it aligns closely with memorized patterns in the training data. However, in the absence of overlap, Template C underperforms due to its rigidity, whereas Template A’s conversational and flexible nature enables better generalization, leading to improved performance.
Facebook OPT Model Results. To assess our approach with a model whose training data has no overlap with Enron, we employ Facebook’s OPT model. Notably, this model’s training dataset excludes not only the Enron dataset but also any email data altogether. The evaluation of the 2.7b parameter version of OPT, as depicted in Figure 5, demonstrates the effectiveness of our proposed method. Similar to the results obtained when training Pythia with Pseudonymized Enron, the model’s performance across various templates remained consistent or improved. Notably, Template A demonstrates a significant performance boost, consistent with the results in Pseudonymized Enron setting, likely due to its low resemblance to email-like data. Additionally, templates B and C were the most challenging due to the same reasons outlined in the preceding paragraph. The observations regarding PEFT methods, as discussed earlier, are also applicable here, highlighting their potential influence on the performance patterns observed across different templates.
3.3.3 Understanding Privacy Risks in Fine-Tuning
Reasons for the Privacy Risks. To find the possible reason for the increased privacy risks, we first show a sample from the generated dataset in Figure 6. The figure illustrates that the structure of the generated data closely resembles that of the test data. Specifically, we found that the total generated data consists of over $60,000$ name-email pairs for around $2000$ generated samples. Such structure similarities may revoke LLMs’ memorization of name and email pairs and lead to privacy risks, as LLMs feed too many name-email pairs during the fine-tuning. Former research verified the observation that repeated sequences can increase memorization in LLMs [22]. In our case, the structural similarity implies repeated sequences of e-mail conversations, such as header lines of former emails including metadata. The structural similarity may strengthen the effect of memorization, as seen in the context of data augmentation [35].
Apart from the structure similarity, we also notice semantic similarity. The semantic similarity score evaluated by the Sentence Transformer  [6] between the generated and the original data is over $0.7$ Figure 6 also reveals that many specific name-email relationships are closely mirrored between the generated data and the test data. For example, the email’s local name can be formed by inserting a "_" or "." between the first name and the second name, or just concatenating the first character in the first name with the last name to form an email address like dhansen for Don Hansen. Such similar relationships present in both the generated data and the original pre-training data can trigger the LLM’s memory, reinforcing its recall of the connections between names and emails.
Learning Rates Impacts on the Privacy Risks. As illustrated in Tirumala et al.’s work [49], the learning rate is a key factor in LLM’s memorization and the model’s final performance. Therefore, we further explore the learning rate’s impacts on both the model’s utility and the extraction success rate in Figure 7. In Figure 7, we observe that the utility and extraction success rates show a similar trend with the learning rate changes. With the increment of the learning rate for fine-tuning, the perplexities after fine-tuning first decrease significantly but will also increase when the learning rate becomes too large. The best learning rate for better utility is around $10^{-6}$ to $5\times 10^{-5}$ for different models. Regarding PII extractions, there are no clear trends across different models concerning varying learning rates. However, one consistent pattern emerges: models fine-tuned with lower learning rates tend to have lower numbers of successful PII extractions. A possible explanation is that slightly larger learning rates enable LLMs to memorize patterns more effectively, leading to higher privacy risks, consistent with the findings in [49]. Therefore, we recommend using the smaller learning for fine-tuning, e.g., around $10^{-6}$, to alleviate the privacy risks while improving the utilities on the target domain.
@@ 3.4 Discussion
In this section, we examine the privacy risks posed to LLMs after fine-tuning them with generated instructional data. Using the Enron email dataset as a case study, we fine-tune Pythia models of varying sizes (410m, 1.4b, 2.8b) with email data generated by a Pythia 12B model. We then assess the privacy risks by performing PII attacks on the Enron dataset, which is related to the pre-training data. The results reveal that, after fine-tuning, the Pythia models are able to extract over $20\%$ more PII data compared to the base model. This finding indicates that fine-tuning with generated data can heighten the model’s privacy risks concerning the pre-training dataset.
To further verify these findings and broaden our evaluation, we utilized Facebook’s OPT model alongside a modified version of the Enron dataset, Psedonymized Enron. The OPT model was chosen because its training data has no overlap with the original Enron dataset or any email data, allowing us to assess privacy risks in a setting lack of direct training correlations. Psedonymized Enron dataset was designed to minimize the overlap with Pythia’s training data, ensuring that any identified risks arose from the model’s learning behavior and not from pre-existing overlaps in the training data. The results show approximately a $40\%$ improvement in PIIs for certain templates. This behavior confirms the findings from experiments where the training and fine-tuning datasets have overlapping content.